# Gravitational Lensing of Gravitational Waves from Astrophysical Sources: Theory, Detection, and Applications

### 07 May 2026 v1

@@ 5.2 Dark Lensed Sirens
The crucial element of this “Dark Lensed Siren” method is identifying and reconstructing the lensed host galaxy, then achieving the sub-arcsecond re-localization of the GW source using GW-measured the time-delay and magnification ratios. There are three necessary key ingredients for the application of this method (see details in Chen et al. 2025a):
@@
Based on the above recipes, Chen et al. (2025a) estimated the constraining power of the “Dark Lensed Siren” on the Hubble constant, and they found that $H_{0}$ can be constrained with $\lesssim 1\%$ precision with a period of about two years detection of the dark lensed events by third-generation GW detectors.
@@ 6 Searching in GWTC-data
Lensed GW events are one type of the main targets for GW detection, which have been searched for in the LVK data from the beginning of the ground-based GW observations. However, there is still no compelling evidence for lensing signatures, i.e., strong lensing with multiple images or diffractive microlensing with frequency-dependent amplification, discovered in the observed GW signals by the LVK network (e.g., Abbott et al. 2021, 2024; The LIGO Scientific Collaboration et al. 2025a), Nevertheless, a number of methods have been proposed and applied to search for such lensed events in real data. Here we summarize the main methods currently applied in searching for lensed GW events.
@@ 6.1 Preliminary Identification of Lensed Pair Candidates
where $p(\Delta t|\mathcal{H_{\rm SL}})$ and $p(\Delta t|\mathcal{H_{\rm U}})$ are the prior probabilities of time-delay $\Delta t$ under the strongly lensed and unlensed hypotheses, respectively. Note that $\mathcal{R}^{\rm gal}$ is dependent on the lensing model and therefore subject to some astrophysical uncertainties. Abbott et al. (2021) selected most promising events pairs by the above quantities, i.e., $\mathcal{B}^{\rm overlap}>50$ (and $\mathcal{R}^{\rm gal}>0.01$ ). In addition, with the rapid development of the machine-learning (ML) technique, it is possible to use ML to classify candidate pairs of lensed events, which is independent of the posterior-overlap analysis (Goyal et al. 2021). By ML approach, one did not need to generate the above Bayesian posterior which may take hours to days, but directly construct the time-frequency and localization sky maps in the order of seconds and thus may be used to conduct preliminary identification efficiently.
@@ 6.2 Joint Parameter Estimation Analysis
Unlike the preliminary identification, the Joint PE analysis operates directly simultaneously on the detector strain data instead of on independently derived posterior samples for each component of the candidate pair. Although computationally intensive, the Joint PE analysis explores the full parameter space and therefore yields substantially stronger tests on the lensing hypothesis than methods that marginalize over or treat parameters separately. There are three estimators in such analysis typically used to quantify the evidence for the strong lensing hypothesis as follows (e.g., Abbott et al. 2021; Lo & Magaña Hernandez 2023; Abbott et al. 2024). (1) Coherence ratio $\mathcal{C_{\rm U}^{\rm L}}$: defined as the ratio of lensed and unlensed evidences, neglecting selection effects and using default priors, which quantifies the confidence of parameters of the GWs aligning with the expectations of lensed GW events, but can not account for the coincident parameter overlap. (2) Population-weighted coherence ratio $\left.\mathcal{C_{\rm U}^{\rm L}}\right|_{\rm pop}$: taking into account the likelihood that GW parameters overlap by chance. However, the value of $\left.\mathcal{C_{\rm U}^{\rm L}}\right|_{\rm pop}$ is largely subject to the choice of both CBC population and lens models. (3) Bayes factor $\mathcal{B}_{\rm U}^{\rm L}$: quantifying the evidence for the strong lensing hypothesis given a detector network and the population model, which explicitly takes into account the selection effects (for example, certain CBC mass ranges and sky orientations are preferentially detected because of the GW detectors’ PSDs). Among these three estimators, the most decisive is the Bayes factor $\mathcal{B}_{\rm U}^{\rm L}$, which is however computationally expensive. A common procedure is to first select pair candidates with high $\mathcal{C_{\rm U}^{\rm L}}$ and high $\mathcal{C_{\rm U}^{\rm L}}\mid_{\rm pop}$ (for example, $\log\mathcal{C_{\rm U}^{\rm L}}>4$), and then compute $\mathcal{B}_{\rm U}^{\rm L}$ only for these candidates.
The above joint PE procedure has been implemented into pipelines, such as GOLUM (Janquart et al. 2023) and HANABI (Lo & Magaña Hernandez 2023). Both the pipelines adopt the nested sampling algorithm Dynesty (Speagle 2020) and implement the joint PE with the help of Bilby (Ashton et al. 2019; Romero-Shaw et al. 2020). In a recent paper by Abbott et al. (2024), HANABI was adopted to estimate $\mathcal{B}_{\rm U}^{\rm L}$ for $17$ pairs, all of which yielded $\mathcal{B}_{\rm U}^{\rm L}<1$. Therefore, it still remains no evidence for lensing signatures in current GWTC data. Notably, Janquart et al. (2023) performed a systematic follow-up analyses of the O3 LVK lensing searches and identified two strong-lensing candidate pairs, i.e., GW191103-GW191105 and GW191230_180458-GW200104_180425, but were unable to obtain high statistical significance for lensing hypothesis.
@@ 6.4 Search for Microlensing Effect
where $\theta$ represents the set of parameters defining an unlensed GW signal, $M_{\rm L}^{z}$ is the redshifted lens mass and $y$ is the dimensionless source position in source plane, and $F(f;M_{\rm L}^{z},y)$ is the amplification factor, which can be estimated by Equation (41). Then we may estimate the Bayes factor $\mathcal{B}_{\rm U}^{\rm Micro}$, i.e., evidence ratio between microlensed and unlensed signal models, for all the events reported in GWTC catalog. Here we note that a high Bayes factor $\mathcal{B}_{\rm U}^{\rm Micro}$ itself does not necessarily mean a conclusive evidence of microlensing in an observed event, since the statistical fluctuation of $\mathcal{B}_{\rm U}^{\rm Micro}$ for unlensed signals may also be high enough and leads to false positives. For example, Abbott et al. (2021) generated unlensed injections with Gaussian noise and recovered them with lensed templates, and they found that the typical statistical fluctuations of the resulting Bayes factors due to false alarm can be as high as $\log\mathcal{B}_{\rm U}^{\rm Micro}\sim 0.75$. Abbott et al. (2024) estimated $\mathcal{B}_{\rm U}^{\rm Micro}$ for all the GW events in O3 and found that the distribution of $\log\mathcal{B}_{\rm U}^{\rm Micro}$ is centered at 0 and the distribution does not deviate significantly with that from unlensed samples. Recently, The LIGO Scientific Collaboration et al. (2025a) searched for the O4a data with similar procedures and found no compelling evidence for the presence of microlensing signatures in current GWTC data.

### 16 May 2026 v2

@@ 5.2 Dark Lensed Sirens
The crucial element of this “Dark Lensed Siren” method is identifying and reconstructing the lensed host galaxy, then achieving the sub-arcsecond re-localization of the GW source using GW-measured the time-delay and magnification ratios. There are three necessary key ingredients for the application of this method (see details in Chen et al. 2025b):
@@
Based on the above recipes, Chen et al. (2025b) estimated the constraining power of the “Dark Lensed Siren” on the Hubble constant, and they found that $H_{0}$ can be constrained with $\lesssim 1\%$ precision with a period of about two years detection of the dark lensed events by third-generation GW detectors.
@@ 6 Searching in GWTC-data
Lensed GW events are one type of the main targets for GW detection, which have been searched for in the LVK data from the beginning of the ground-based GW observations. One motivation is that lensing magnification can make intrinsically lower-mass, higher-redshift GW events appear as heavier systems, leading to the proposal that some high-mass BBHs observed by the LVK could be lensed events (Diego et al. 2021). Although this possibility illustrates the importance of dedicated lensing analyses, recent studies find that lensing magnification cannot consistently explain the observed high-mass BBH population (Harshe et al. 2026). More generally, there is still no compelling evidence for lensing signatures, i.e., strong lensing with multiple images or diffractive microlensing with frequency-dependent amplification, in the GW signals observed by the LVK network (e.g., Abbott et al. 2021, 2024; The LIGO Scientific Collaboration et al. 2025e). Nevertheless, a number of methods have been proposed and applied to search for such lensed events in real data. Here we summarize the main methods currently applied in searching for lensed GW events.
@@ 6.1 Preliminary Identification of Lensed Pair Candidates
where $p(\Delta t|\mathcal{H_{\rm SL}})$ and $p(\Delta t|\mathcal{H_{\rm U}})$ are the prior probabilities of time-delay $\Delta t$ under the strongly lensed and unlensed hypotheses, respectively. Note that $\mathcal{R}^{\rm gal}$ is dependent on the lensing model and therefore subject to some astrophysical uncertainties. Abbott et al. (2021) selected most promising events pairs by the above quantities, i.e., $\mathcal{B}^{\rm overlap}>50$ (and $\mathcal{R}^{\rm gal}>0.01$ ). In addition, with the rapid development of the machine-learning (ML) technique, it is possible to use ML to classify candidate pairs of lensed events, which is independent of the posterior-overlap analysis (Goyal et al. 2021). By ML approach, one did not need to generate the above Bayesian posterior which may take hours to days, but directly construct the time-frequency and localization sky maps in the order of seconds and thus may be used to conduct preliminary identification efficiently. Note here that we also comment that recently Barsode et al. (2025b) developed a modified version of the posterior overlap method and find it is possible to identify $\sim 65\%$ lensed events with LVK efficiently.
@@ 6.2 Joint Parameter Estimation Analysis
Unlike the preliminary identification, the Joint PE analysis operates directly simultaneously on the detector strain data instead of on independently derived posterior samples for each component of the candidate pair. Although computationally intensive, the Joint PE analysis explores the full parameter space and therefore yields substantially stronger tests on the lensing hypothesis than methods that marginalize over or treat parameters separately. There are three estimators in such analysis typically used to quantify the evidence for the strong lensing hypothesis as follows (e.g., Abbott et al. 2021; Lo and Magaña Hernandez 2023; Abbott et al. 2024). (1) Coherence ratio $\mathcal{C_{\rm U}^{\rm L}}$: defined as the ratio of lensed and unlensed evidences, neglecting selection effects and using default priors, which quantifies the confidence of parameters of the GWs aligning with the expectations of lensed GW events, but can not account for the coincident parameter overlap. (2) Population-weighted coherence ratio $\left.\mathcal{C_{\rm U}^{\rm L}}\right|_{\rm pop}$: taking into account the likelihood that GW parameters overlap by chance. However, the value of $\left.\mathcal{C_{\rm U}^{\rm L}}\right|_{\rm pop}$ is largely subject to the choice of both CBC population and lens models. (3) Bayes factor $\mathcal{B}_{\rm U}^{\rm L}$: quantifying the evidence for the strong lensing hypothesis given a detector network and the population model, which explicitly takes into account the selection effects (for example, certain CBC mass ranges and sky orientations are preferentially detected because of the GW detectors’ PSDs). Among these three estimators, the most decisive is the Bayes factor $\mathcal{B}_{\rm U}^{\rm L}$, which is however computationally expensive. A common procedure is to first select pair candidates with high $\mathcal{C_{\rm U}^{\rm L}}$ and high $\mathcal{C_{\rm U}^{\rm L}}\mid_{\rm pop}$ (for example, $\log\mathcal{C_{\rm U}^{\rm L}}>4$), and then compute $\mathcal{B}_{\rm U}^{\rm L}$ only for these candidates.
The above joint PE procedure has been implemented into pipelines, such as GOLUM (Janquart et al. 2023) and HANABI (Lo and Magaña Hernandez 2023). Both the pipelines adopt the nested sampling algorithm Dynesty (Speagle 2020) and implement the joint PE with the help of Bilby (Ashton et al. 2019; Romero-Shaw et al. 2020). In a recent paper by Abbott et al. (2024), HANABI was adopted to estimate $\mathcal{B}_{\rm U}^{\rm L}$ for $17$ pairs, all of which yielded $\mathcal{B}_{\rm U}^{\rm L}<1$. Therefore, it still remains no evidence for lensing signatures in current GWTC data. Notably, Janquart et al. (2023) performed a systematic follow-up analyses of the O3 LVK lensing searches and identified two strong-lensing candidate pairs, i.e., GW191103-GW191105 and GW191230_180458-GW200104_180425, but were unable to obtain high statistical significance for lensing hypothesis.
@@ 6.4 Search for Microlensing Effect
where $\theta$ represents the set of parameters defining an unlensed GW signal, $M_{\rm L}^{z}$ is the redshifted lens mass and $y$ is the dimensionless source position in source plane, and $F(f;M_{\rm L}^{z},y)$ is the amplification factor, which can be estimated by Equation (41). Then we may estimate the Bayes factor $\mathcal{B}_{\rm U}^{\rm Micro}$, i.e., evidence ratio between microlensed and unlensed signal models, for all the events reported in GWTC catalog. Here we note that a high Bayes factor $\mathcal{B}_{\rm U}^{\rm Micro}$ itself does not necessarily mean a conclusive evidence of microlensing in an observed event, since the statistical fluctuation of $\mathcal{B}_{\rm U}^{\rm Micro}$ for unlensed signals may also be high enough and leads to false positives. For example, Abbott et al. (2021) generated unlensed injections with Gaussian noise and recovered them with lensed templates, and they found that the typical statistical fluctuations of the resulting Bayes factors due to false alarm can be as high as $\log\mathcal{B}_{\rm U}^{\rm Micro}\sim 0.75$. Abbott et al. (2024) estimated $\mathcal{B}_{\rm U}^{\rm Micro}$ for all the GW events in O3 and found that the distribution of $\log\mathcal{B}_{\rm U}^{\rm Micro}$ is centered at 0 and the distribution does not deviate significantly with that from unlensed samples. Recently, The LIGO Scientific Collaboration et al. (2025e) searched for the O4a data with similar procedures and found no compelling evidence for the presence of microlensing signatures in current GWTC data.